OpenStudy (anonymous):

Multiply and simplify: (3 + √3) (5 - 4√3) Answers: 3 15-4√3 3-7√3

OpenStudy (debbieg):



OpenStudy (debbieg):

Just FOIL type multiplication. Just like with 2 binomials. Then simplify each term, combine the like terms.

OpenStudy (anonymous):

wouldn't it be B? O_o

OpenStudy (anonymous):

actually no, I think its A according to the foil

OpenStudy (debbieg):

What do you get for the 4 terms of the FOIL?

OpenStudy (anonymous):

am i right?

OpenStudy (debbieg):

Don't worry about the answer choices. What do you get for the 4 terms of the FOIL?

OpenStudy (anonymous):

wait i multiply all of them?

OpenStudy (debbieg):

it's just like foil: (a+b)(c+d)=ac+ad+bc+bd

OpenStudy (anonymous):

okay, now I'm totally lost :/

OpenStudy (debbieg):

why? do you know what "FOIL" is? how to multiply 2 binomials?

OpenStudy (anonymous):

i guess not

OpenStudy (debbieg):

If you prefer, you can think of it as distribution. Same thing: (a+b)(c+d)=a(c+d) + b(c+d) =ac +ad +bc +bd That works, right? you know how to distribute?

OpenStudy (debbieg):

Do it that way.

OpenStudy (debbieg):

@sara17 stop giving answers please. OpenStudy values the Learning process - not the ‘Give you an answer’ process. Don’t post only answers - guide the asker to a solution. http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

ok.. so sorry

OpenStudy (anonymous):

no idea how to do any of this

OpenStudy (anonymous):

i forgot everything

OpenStudy (debbieg):

Are you reading what I wrote above? Use distribution: (3 + √3) (5 - 4√3)= 3(5 - 4√3) + √3(5 - 4√3) right? now what?

OpenStudy (anonymous):

O_o

OpenStudy (debbieg):

If you forgot things, then you need to do some extra work to re-learn them. Not just give up.

OpenStudy (anonymous):

iknow i know i just that right now i only have a few days to finish my online course and i really want to pass

OpenStudy (anonymous):

so whats the next step?

OpenStudy (debbieg):

If you had: 3(5 - x), what would you do? You would distribute, right? 3(5 - x)=3(5) - 3x Then, for 3(5 - 4√3) you do the same thing: 3(5 - 4√3) =3(5)-3(4√3) That's the first half of the distribution. Do the same for the other part.

OpenStudy (anonymous):

for this (3 + √3) ??

OpenStudy (debbieg):

I don't see that in the problem anywhere... I meant for this: √3(5 - 4√3)

OpenStudy (debbieg):

you have this: (3 + √3) (5 - 4√3) I showed you, you distribute like this: (3 + √3) (5 - 4√3) = 3(5 - 4√3) + √3(5 - 4√3) Now you distribute again, through each of those parens.... I did the first one: (3 + √3) (5 - 4√3) = 3(5 - 4√3) + √3(5 - 4√3) =3(5)-3(4√3)+ √3(5 - 4√3) ^^^^^^^^^ Now you do the 2nd one ^^

OpenStudy (anonymous):

i don't even know what the second one is! :/ forget it -.-

OpenStudy (anonymous):

thx for trying